OpenStudy (anonymous):

If you roll two dice,what is the probability that you will roll 5 each?

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

@GeniousCreation

OpenStudy (kendricklamar2014):

Probability of this happening: (1/6)(1/6)

OpenStudy (kendricklamar2014):

Total probability: 4/36

OpenStudy (kendricklamar2014):

\[\frac{ 4 }{ 36 } = \frac{ 1 }{ 9 }\] \[Answer = \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

thankyou

OpenStudy (kendricklamar2014):

You're Welcome